Circular No. 7506

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 T1 (SOHO) Further to IAUC 7445 and 7480, D. Biesecker reports measurements

for a Kreutz sungrazing comet found by J. Danaher in LASCO C3 images at the SOHO website. Biesecker provides some V magnitudes (estimated uncertainties 0.1 mag): Oct. 10.096 UT, 7.1; 10.221, 6.6; 10.263, 6.4; 10.321, 6.3. The reduced positions and parabolic orbit by B. G. Marsden are given on MPEC 2000-T33.

2000 UT  α_{2000}  δ_{2000}
Oct. 9.821  12^{h}29.^{m}0  − 8^{o}45′

SUPERNOVA 2000dr IN IC 1610 N. B. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), re-

ports: “R. C. Smith obtained a spectrum (range 480–680 nm, resolution 1000) of SN 2000dr (cf. IAUC 7505) on Oct. 8.1 UT with the CTIO 1.5m telescope (+ facility spectrograph). The spectrum shows that this is a type-Ia supernova, based on the presence of strong Si II (597, 635 nm) absorption. The high photospheric velocities for these lines (12 500 km/s) and the lack of developed S II lines at 547 and 563 nm indicate that this supernova is at least a week before maximum.”

J1859+226 C. Sanchez-Fernandez, Laboratorio de Astrofísica Espacial y Física

Fundamental; C. Zurita, J. Casares, and T. Shahbaz, Instituto de Astrofisica de Canarias; and A. Castro-Tirado, Instituto de Astrofisica de Andalucia, write: “Continuous R-band monitoring obtained since Aug. 24 with the IAC80 telescope at the Observatorio del Teide indicates that J1859+226 has reached quiescence at R ∼ 23, consistent with the upper limit reported on IAUC 7494. R-band CCD images taken in good seeing with the 4.2-m William Herschel Telescope at La Palma on Sept. 27–28 revealed the presence of a faint star (R = 23.05 ± 0.04) located 1′′.4 north of the target. J1859+226 is clearly detected at R = 22.90 ± 0.03 and exhibits a sinusoidal modulation, with a period of 0.19 ± 0.01 day and a peakto-peak amplitude of ∼ 0.4 mag. Assuming a double-humped ellipsoidal lightcurve, as is typical of quiescent soft-x-ray transients, we estimate an orbital period of 0.38 ± 0.01 day. This value is in good agreement with the period reported on IAUC 7388, where x-ray reprocessing in the disk and the companion star was the dominant source of light.”

2000 October 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green